Nine people are going on a skiing trip in 3 cars
that hold 2, 4, and 5 passengers, respectively. In how
many ways is it possible to transport the 9 people to
the ski lodge, using all cars?

Answer by Alan3354(69443)

Nine people are going on a skiing trip in 3 cars
that hold 2, 4, and 5 passengers, respectively. In how
many ways is it possible to transport the 9 people to
the ski lodge, using all cars?
====================
2+2+5 = 9
2+3+4 = 9
2+4+3 = 9
1+4+4 = 9
1+3+5 = 9